OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b?

OpenStudy (mokeira):

\[a-b=21\] \[5a-2b=18\] we can use substitution \[a-b=21\] \[a=21+b\] Substitute in the second equation \[5(21+b)-2b=18\] \[105+5b-2b=18\] \[3b=-87\] \[b=-29\] Go back to first equation \[a-b=21\] \[a--29=21\] \[a+29=21\] \[a=-8\]

OpenStudy (triciaal):

given a - b = 21 given 5 a - 2 b = 18 solve simultaneously by substitution or by elimination to use substitution: from the first equation let a = 21 + b then 5(21 + b ) - 2 b = 18 105 + 5 b - 2 b = 18 b = -87/3 = -29 a = 21- 29 = -8

OpenStudy (triciaal):

to use elimination multiply the first equation by -2 add the second -2 a + 2 b = -42 5 a -2 a + 2 b - 2 b = -42 + 18 3 a = -24 a = -8 substitute for a in original equation to get b